### v1

# Patterns of X-ray and -ray Flares: from Blazar to Maganetar and Sun

Abstract
Using Gaussian process methods, we analyzed the light curves of three extreme solar X-ray flares observed by the RHESSI satellite. Their variability characteristics were then compared with those of HXMT-HE X-ray burst (XRB; in SGR 1935+2154) associated with fast radio burst (FRB) 200428 and blazar -ray giant flares, to investigate the origins of these extreme flaring events. The variability patterns of the solar X-ray flares follow the stochastically driven damped simple harmonic oscillator (SHO) model. The derived timescales and (corresponding to PSD breaks) are in the range of 4-7 s and 16-53 s, respectively. The FRB-associated HXMT-HE burst has a value near 0.3, matching those of the solar flares occurred on 23 July 2002 (flare 1) and 3 November 2003 (flare 2). By contrast, blazar -ray giant flares show Q 0.3, similar to the solar flare occurred on 25 February 2014 (flare 3). We proposed that the critically damped state of the system may be the condition triggering the association between the XRB in SGR 1935+2154 and the FRB. In this scenario, the critical damping value of the system is around 0.3, not the theoretical 0.5. The similarity in values might imply that the FRB-associated HXMT-HE XRB and solar X-ray flares 1 2 share comparable dynamic behavior, while blazar -ray flares and solar X-ray flare 3 exhibit another distinct but similar dynamic behavior. Like solar X-ray flares, these extreme flares may all be related to the magnetic reconnection process.
UTF8gbsn
1 Introduction
Flaring phenomena in different astronomical objects, such as solar flares, magnetar X-ray bursts (XRBs), -ray bursts and large flares from active galactic nuclei (AGNs), typically involve the processes of energy release, particle acceleration, and magnetic activity, with their physical environments, underlying mechanisms and radiation signatures potentially differing or exhibiting similarities in certain aspects. Studying the time-domain characteristics of different flare events allows us to gain deeper insights into involved physical processes in extreme astrophysical conditions.
Blazars are a distinct subclass of AGNs, characterized by their relativistic jets pointing directly toward the Earth (e.g., Ackermann etĀ al., 2015). They emit radiation across entire electromagnetic spectrum. Blazars typically exhibit strong variability, which results in prominent time-domain characteristics in multi-wavelength observations. The long-term variability of blazars has been studied through Gaussian Process method (GP; Zhang etĀ al., 2021, 2022, 2023). The main findings indicate that such variability follows the damped random walk (DRW) model, with the characteristic timescales associated with the thermal instability timescale of the accretion disk (Zhang etĀ al., 2022, 2023). As prominent blazars, both 3C 454.3 and 3C 279 exhibit -ray flaring events whose light curves were suggested to follow the stochastically driven damped simple harmonic oscillator (SHO) model in the overdamped regime (Zhang etĀ al., 2025b). This variability behavior is typically attributed to internal processes within the innermost regions of their relativistic jets (e.g., Ackermann etĀ al., 2016; Shukla etĀ al., 2018).
Magnetars are a type of neutron stars with extremely strong magnetic fields ( G; see Kaspi & Beloborodov (2017) for a review), whose intense magnetic activity can trigger multi-band radiation bursts, e.g., XRBs, radio bursts (may include fast radio bursts, FRBs), and -ray bursts (Turolla etĀ al., 2015). Studies suggest that the energy release from magnetar bursts could be linked to direct magnetic reconnection in the magnetosphere, driven by the intense magnetic field (e.g., Beloborodov, 2013; Mereghetti etĀ al., 2015). Alternatively, these flares might be triggered by shear waves released during crustal rupture, which transmit energy to the magnetosphere through Alfvn waves, initiating secondary magnetic reconnection (e.g., Li & Beloborodov, 2015; Yuan etĀ al., 2020). SGR J1935+2154 is a Galactic magnetar, of which hundreds of XRBs were recorded by different detectors, e.g., Insight-HXMT (Li etĀ al., 2021), Neutron Star Interior Composition Explorer (NICER; Yuan etĀ al., 2020). A non-thermal XRB of SGR J1935+2154 on 2020 April 28th has been discovered to be accompanied by emission of the FRB 200428 (CHIME/FRB Collaboration etĀ al., 2020; Li etĀ al., 2021), suggesting the non-thermal XRB and FRB 200428 share the same physical origin in an explosive event from SGR J1935+2154. For XRBs in SGR J1935+2154, Tang etĀ al. (2024) used GP method to characterize their variability. They found that both normal XRBs and FRB-associated HXMT XRB generally show variability consistent with the SHO patern, the derived timescales can put constraints on the production mechanisms of XRBs.
Solar flares are explosive phenomena in the solar atmosphere, which can emit electromagnetic emission extending to -rays within a short time. Extensive research has been conducted on particle acceleration, radiation mechanisms, and emission regions in solar flares (e.g., Holman etĀ al., 2011; Chen etĀ al., 2020; Share etĀ al., 2024). It is generally agreed that energy release is usually attributed to magnetic reconnection occurring on the current sheet (e.g., Chen, 2011; Benz, 2017; Share etĀ al., 2018; Chen etĀ al., 2020). Research indicates that the light curves of hard X-rays ( keV) display fast pulsating structures with periods ranging from milliseconds to seconds, closely related to the acceleration of high-energy electrons induced by magnetic reconnection, with non-thermal bremsstrahlung as the primary radiation mechanism (e.g., Holman etĀ al., 2011; Kontar etĀ al., 2019; Krucker etĀ al., 2020).
Building on the analysis of the variability patterns of blazar -ray flares and magnetar XRBs, we would like to explore the variability of the solar flares by GP method in this work. Furthermore, we will compare the burst properties of solar X-ray flares, magnetar XRBs, and -ray blazar outbursts, focusing on investigating the specific mechanisms driving high-energy flare phenomena across these different types of sources. The structure of this paper is as follows: Section 2 describes the construction of the hard X-ray light curves of solar flares, as well as the models and methods used to simulate the light curves. Section 3 displays the fitting results for solar X-ray flares by three different kernels, and compares the variability characteristics of solar X-ray flares, magnetar XRBs, and -ray outbursts of blazars. In Section 4, we discuss the possible origins of the three types of bursts and the potential issues associated with them, based on the results.
2 Method and data processing
The Reuven Ramaty High Energy Solar Spectroscopic Imager (RHESSI) is a NASA Small Explorer Mission Lin etĀ al. (2002). It is capable of performing both imaging and spectroscopic analyses across a broad energy spectrum, ranging from soft X-rays to -rays (3 keV to 17 MeV). The highest energy resolution is as fine as 1 keV. In this study, we utilize the SSWIDL software to extract RHESSI hard X-ray lightcurves, with the temporal resolution of flare light curves set at 4 seconds.
We perform a detailed analysis of the flare light curves using the GP method (Aigrain & Foreman-Mackey, 2023). The mean function of the GP is chosen to be a constant equal to the average count rate of the light curve for each flare. The covariance (kernel) function of the GP therefore serves to represent all variability characteristics of the flares, with its selection being a critical issue in GP. Following empirical practice, we opt for three kernel functions to fit the light curves, i.e., DRW, SHO, and Matrn.
DRW is typically used to capture random fluctuations that gradually decay, with its PSD performing a broken power-law form (e.g., MacLeod etĀ al., 2010; Burke etĀ al., 2021). SHO is most commonly used to describe periodic behaviors in time series, corresponding to the underdamped regime (with the quality factor of the oscillator ), where its PSD shows a peak (e.g., Yang etĀ al., 2021; Zhang etĀ al., 2021). In the overdamped regime (), the PSD is more complex, with the low-frequency part matching the DRW PSD, where the spectral index transitions from -2 to 0 111The change of spectrum index mentioned in this paper refers to the change from high frequencies to low frequencies. (characteristic timescale denoted as ). At high frequencies, the spectral index drops to -4 (the timescale corresponding to the transition from spectral index -4 to -2). In the critically damped regime (), the PSD tends to exhibit a direct transition from a spectral index of -4 to 0 (Kasliwal etĀ al., 2017; Moreno etĀ al., 2019). Different regimes may describe distinct characteristics of the data. The Matrn kernel can capture data structures with large fluctuations, and its PSD features a direct transition from a spectral index of -4 to 0 (Rasmussen & Williams, 2006).
For Matrn, the function is as follow:
which is derived by setting the smoothness parameter (half-integer) of the Matrn class kernels to 3/2. and represent the characteristic timescale and the amplitude respectively. The details of DRW and SHO models can refer to our previous works (Zhang etĀ al., 2021, 2022, 2023). Here, we supplement two timescale (broken timescale of PSD) formulas for the SHO model in overdamped modeļ¼
we use a tool of Celerite (Foreman-Mackey etĀ al., 2017) to perform GP fitting time series. Markov chain Monte Carlo (MCMC) technology is used when implementing the light curve fitting. We implement it through the package emcee 222https://github.com/dfm/emcee (Foreman-Mackey etĀ al., 2013), taking the last 30,000 steps out of 50,000 total MCMC steps as effective samples. Actually, before running MCMC, we choose the initial parameter values through calculating the maximum likelihood. Then, we can construct the final PSD by taking the median of the 30,000 PSD values corresponding to each frequency, and plot 30,000 groups of parameters as a distribution, i.e., a posterior distribution of the model parameters.
3 results
3.1 results of the solar X-ray flares
Three bright solar hard X-ray flares, occurring on 23 July 2002 (flare 1), 3 November 2003 (flare 2) and 25 February 2014 (flare 3) in the 50-100 keV and 100-300 keV energies, are analyzed, with data obtained from the RHESSI satellite. The specific time periods for the flares are listed in Table 1. DRW, SHO and Matrn models are chosen to describe these light curves. Considering the fit between the observed data and modeled light curve, the distribution of standard residuals, as well as the autocorrelation functions (ACF) of both the standard residuals and their squares (the goodness-of-fit criteria described in detail in Zhang etĀ al. (2022)), we observe no significant differences between the three models for any of the flares. Further evaluation using the corrected Akaike information criterion (), with values listed in TableĀ 1, shows that the SHO model outperforms the DRW and Matrn models with for all flares across both energy bands.
We present the SHO fitting plots for the solar flares in the 50-100 keV and 100-300 keV energy bands, as shown in FigureĀ 1 and FigureĀ 2. From left to right, the panels display the results for flare 1, flare 2 and flare 3 respectively. Although the residual distribution for the flare 3 does not fit a normal distribution well, the concentration of residuals around zero and the lack of trends or patterns suggest that the SHO model is still considered valid.
The corresponding posterior probability density distributions of the SHO parameters are shown in FigureĀ 3. The SHO parameters for all flares are well constrained, with the specific values provided in TableĀ 1.
Now we focus on the PSDs shown in FigureĀ 4. Flare 1 and flare 3 have PSD indexes changing from -4 to -2 and then to 0. That is, the timescales of and are constrained. The two timescales derived from flare 3 in 100-300 keV are in close proximity to each other. However, the PSD indexes of flare 2 in two energy bands evolve from -4 to -2, with only constrained. The spectral indexes of PSDs and the characteristic timescales of these solar flares are listed in the TableĀ 2. and of the solar flares are in the range of 4-7 s and 16-53 s, respectively. The relationship between these characteristic timescales and the broken frequency of the PSD is .
Solar flare | Time period | Parameter of SHO | |||||
ln | ln Q | ln | SHO | DRW | Matrn | ||
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
flare 1 | 2002-07-23 00:27:18-00:41:10 | 2772 | 3461 | 3372 | |||
flare 2 | 2003-11-03 01:15:34-01:25:30 | 1775 | 1987 | 2321 | |||
flare 3 | 2014-02-25 00:43:22-00:50:02 | 1346 | 1658 | 1601 | |||
flare 1 | 2002-07-23 00:27:22-00:41:10 | 2237 | 2352 | 2860 | |||
flare 2 | 2003-11-03 01:15:22-00:25:14 | 1758 | 1959 | 2319 | |||
flare 3 | 2014-02-25 00:43:22-00:50:02 | 1235 | 1414 | 1483 |
Note. ā (1) source name, (2) time period of the flare, (3)(4)(5) posterior parameters of modeling light curves with SHO model, (6)(7)(8) values of SHO, DRW and Matrn model. The uncertainties of model parameters represent confidence intervals. The upper part shows the results for 50-100 keV, and the lower part for 100-300 keV.
3.2 Comparison of flares from blazars, magnetars and the Sun
In previous work, we have analyzed the -ray blazar flares (of 3C 454.3 and 3C 279) and the XRBs occurred in magnetar SGR 1935+2154 by GP. The light curves of these flares also follow the SHO model. By using the solar X-ray flare results as a reference, we seek to compare the variability patterns and timescales of blazar -ray flares and XRBs in magnetars, further investigating the physical origins of these distinct flare events.
Considering that the exact mode of the SHO model is determined by the quality oscillation factor , we establish a two-dimensional relationship between the peak luminosity-to-mass ratio of the flares and the parameter derived from the SHO model fits, as shown in FigureĀ 5. For solar flares (the green part in FigureĀ Ā 5), the peak luminosity estimated as erg/s is used. The values of flare 1 and flare 2 are generally near or below 0.3, while flare 3 appears to have values significantly larger than 0.3 and develops to larger values (near or larger than 0.5).
There is a unique XRB in SGR 1935+2154 that is associated with FRB 200428. Its high-energy (HE) data from Insight-HXMT show the best correspondence with the FRB radio data. The value of this XRB in HE is approximately 0.3, which is comparable to that of solar X-ray flares 1 2, yet its peak luminosity-to-mass ratio exceeds that of solar flares by roughly 8-10 orders of magnitude (see FigureĀ 5). The normal XRBs in SGR 1935+2154 possess values spanning a wide range (Tang etĀ al., 2024), without displaying in the figure.
For -ray blazar flares (the blue part), there is a difference of 10-11 orders of magnitude for peak luminosity-to-mass ratio compared to solar X-ray flares. They have values generally larger than 0.3, which is similar to that of solar flare 3. We add the results of -ray long-term variability of 3C 454.3 and 3C 279 (the baby blue part) as a comparison of the blazar flares. Both their values () and the peak luminosity-to-mass ratio are lower than those of the flares. Blazar PG 1553+113 and non-blazar PKS 0521-36 (the orange arrows) exhibit long-term -ray variability with a quasi-periodic oscillation (QPO) signal, and their values deviate significantly from the others shown in FigureĀ 5. It should be noted that the long-term -ray variability of PG 1553+113 is not dominated by this periodic component, but rather by a component described by the Matrn model (Zhang etĀ al., 2021, 2025a).
Additionally, we provide a summary of the PSD characteristics and variability timescales for the analyzed flares in TableĀ 2. The FRB-associated HXMT-HE XRB, like solar flares, possesses a PSD structure with an index changing from -4 to -2, but the timescale of the HXMT-HE XRB is significantly smaller. As a result of data constraints, both the FRB-associated HXMT-HE XRB and solar flare 2 are missing a PSD structure with an index changing from -2 to 0. The blazar -ray flares have the longest observed characteristic timescales, with PSD indexes changing from -4 to 0.
We further simulate the PSDs of SHO for different values and show them in FigureĀ 6. At low values of , the PSD displays two distinct breaks. With an increase in , the two bending points gradually approach each other, merging into almost one break at =0.5. When , the peak feature (QPO signal) becomes visible. This can serve as a reference for the results we have fitted.
Source | Type of PSD | Timescale | Timescale |
the solar X-ray flare 1 | -4 -2 0 | 4-5 s | 43-53 s |
the solar X-ray flare 3 | -4 -2 0 | 5-6 s | 16-23 s |
the solar X-ray flare 2 | -4 -2 | 4-7 s | |
the blazar -ray flares | -4 0 | 2-5 days | 2-5 days |
HXMT-HE burst associated with FRB 200428 | -4 -2 | 0.005 s |
4 Discussion
We have analyzed three bright solar X-ray flares that occurred on 23 July 2002, 3 November 2003 and 25 February 2014 using the GP method. We fitted the light curves of these flares in the 50-100 keV and 100-300 keV energy bands with the SHO, DRW and Matrn models. The results indicate that the solar X-ray flares prefer the SHO model. Both flare 1 and flare 3 constrain two timescales and (PSD indexes: -4 -2 0), with the two timescales of flare 3 being relatively close to each other. While for flare 2, due to the limitation of data length, only is constrained (PSD indexes: -4 -2). for the solar X-ray flares are in the range of 4-7 s, and are in the range of 16-53 s.
The variability of the solar X-ray flares, the blazar -ray flares and the FRB-associated HXMT-HE XRB in SGR 1935+2154 all follow SHO model, yet they exhibit differences in values, PSDs, and characteristic timescales as shown in FigureĀ 5 and TableĀ 2.
We aim to investigate the underlying mechanisms of blazar extreme -ray flares and XRBs in magnetar through a comparative framework, leveraging insights from solar X-ray flares fitting results and their associated physical origins. The driving mechanism of solar X-ray flares is widely attributed to magnetic reconnection, which converts stored magnetic energy into plasma heating and particle acceleration (Krucker etĀ al., 2008; Aschwanden, 2019). In the solar X-ray flare system and in the framework of magnetic reconnection, the timescale (4-7 s) likely corresponds to the energy injection phase, and the flat decay timescale (16-53 s) may reflect subsequent dissipative processes, such as plasma cooling and particle thermalization (e.g., Bian etĀ al., 2012; Aschwanden etĀ al., 2015). The specific dynamical behavior is intrinsically governed by the -factor of the system. Solar X-ray flare 3 exhibits different values compared to flares 1 and 2, potentially due to differences in magnetic topology or plasma environment (Benz, 2017).
The discovery of the FRB-associated HXMT-HE XRB in SGR 1935+2154 suggests that a physical process in magnetar activity (such as magnetic reconnection or oscillation modes) simultaneously leads to the generation of both XRBs and FRBs (CHIME/FRB Collaboration etĀ al., 2020; Li etĀ al., 2021). The critically damped state of a system could be the condition triggering this combined event. In this case, the critically damped regime of this actual system corresponds to , which differs from the theoretical value of . In the comparative framework for the different types of extreme flares, the FRB-associated HXMT-HE XRB have similar values to that of solar flares 1 2, indicating similar dynamic behavior. The difference in characteristic timescale may be attributed to the difference in magnetic field strength. In addition, the ultra-short timescales of FRB-associated HXMT-HE XRB in SGR 1935+2154 (0.005 s) are consistent with Alfvnic crossing times in magnetar magnetospheres (Li & Beloborodov, 2015; Yuan etĀ al., 2020), supporting fast magnetic reconnection triggered by sudden magnetic field reconfiguration.
For blazar -ray flares, the central values of exhibit an upward shift from 0.3, similar to that of solar X-ray flare 3, indicating that they may share certain common dynamic characteristics within their respective systems. A 13-hour quasi-periodic oscillation (QPO) signal detected across -ray flux, optical flux, and linear polarization during the flaring state of BL Lacertae has been reported by Jorstad etĀ al. (2022), in which case the values could lie in the range greater than 1. This result strongly supports an outburst mechanism, i.e., current-driven kink instabilities (e.g., Dong etĀ al., 2020), that modulates the reconnection process.
The diversity in dynamical behaviors among solar X-ray flares, magnetar XRBs, and blazar flares likely stems from differences in magnetic field geometries (e.g., solar loops vs. relativistic jets) and particle acceleration efficiency. For instance, the 2ā5 day timescales of blazar -ray flares may correspond to the light-crossing time of compact emission regions, where it may be the magnetic reconnection in turbulent jets driving rapid particle acceleration. While the second-scale solar X-ray flares and millisecond-scale XRBs in magnetars may reflect energy release processes directly. These high-energy outburst events may share a fundamental reliance on magnetic energy dissipation, such as magnetic reconnection.
An ongoing issue is that, in the light curve fitting, although the ACF of the residual squares mostly falls within the white noise range, it does not randomly distribute around zero, particularly for the solar X-ray flare 3. Future research requires more accurate models and higher-quality data.